Obtain the amount of Co necessary to provide a radioactive source of 8.0 Ci strength. The half-life of Co is 5.3 years?

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,

lamda= ln2/half life
1 Ci=3.7 × 10^10 decays per second
half life=5.3*365*24*60*60=1.671408*10^8
dN/dt=lamda*N=8*3.7 × 10^10
=>N=8*3.7 × 10^10/lamda=8*3.7 × 10^10*1.671408*10^8/ln2 moles

Thanks & Regards
Arun Kumar
Askiitians Faculty
IIT Delhi